# Landing with the Score: Riemannian Optimization through Denoising

### v1

@@ Appendix C Variance-exploding diffusion models
Appendix D Proof of Theorem 1
D.1 Nonasymptotic Laplace method
@@
for some functions $f,h:\mathcal{V}\to\mathbb{R}$ from an open set $\mathcal{V}\subseteq\mathbb{R}^{k}$, where $h$ is non-negative and attains a unique minimum in, say, $z=0\in\mathcal{V}$. While there have been results providing such an error estimation to the first order expansion Inglot & Majerski (2014); Majerski (2015); Lapinski (2019), we need an error estimate up to the second order expansion, as provided in Theorem 14. For the sake of completeness we also include the statement and proof for the first order expansion in Theorem 13. Before we state the results, we need to introduce some notation and conventions. For any $f\in C^{0}(\mathcal{V})$ and $h\in C^{2}(\mathcal{V})$, respectively we write
@@
D.2 Local version of Theorem 1
@@
D.3 Bounds in terms of distribution and manifold parameters
@@ D.3.1 Conditions (24) and (26) for $h$
D.3.2 Lower bound on $\lambda_{\min}(\Sigma)$ and upper bound on $\sqrt{\det\Sigma}$
@@ D.3.3 Condition (24) for $\overline{f}_{2}^{jl}$
D.3.4 Upper bound on $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}$, $\lVert f_{1}^{j}\rVert_{L^{1}(\mathcal{V})}$ and $\lVert\overline{f}_{2}^{jl}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}$ and $\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert$
@@
D.4 Proof of Theorem 1
We apply Theorem 15 to every point $x\in\mathcal{T}(\tau)$ and $p=\pi(x)$ with the graph chart $\psi=\psi_{p}$. It remains to provide universal constants for the bounds in Theorem 15 (i) and (ii). First we bound $E_{1}(\sigma;f_{0},h,r,\theta)$ with $E_{1}$ given in (31). By Section D.3 we can take $r_{\tau_{\mathcal{M}},M}=\eta=\min\{\tau_{\mathcal{M}},M\}/4$, $c_{\tau_{\mathcal{M}},M}=\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}$ and $C_{\tau_{\mathcal{M}},M,\mu}=\max\{C_{\tau_{\mathcal{M}},M}^{(1)},C_{\tau_{\mathcal{M}},M,\mu}^{(2)}\}$ and hence, due to $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}\leq 1$,
@@ D.4 Proof of Theorem 1
D.5 A useful corollary to Theorem 1
@@
Suppose that for some $s:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\tau\in(0,\tau_{\mathcal{M}})$ and $\epsilon\leq\tau$ we have
$\lVert s(x)-\pi(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (42)
Then $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$. Moreover, for $x\in\mathcal{T}(\epsilon)$ we have $\lVert s(x)-x\rVert\leq 2\epsilon$ and for $x\in\mathcal{T}(\tau)$ we have $\lVert s(x)-x\rVert\leq 2\tau$.
@@
From (42) and the fact that $\pi(x)\in\mathcal{M}$ we clearly have $s(x)\in\mathcal{T}(\epsilon)$ for any $x\in\mathcal{T}(\tau)$, i.e. $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$. The last claims follows then from the triangle inequality $\lVert s(x)-x\rVert\leq\lVert s(x)-\pi(x)\rVert+\lVert\pi(x)-x\rVert$. ∎
Appendix E Approximate Riemannian gradient flow with landing
In this section $C=\lVert\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})})$ denote the supremum of $\nabla f$ and the Lipschitz constant of $\nabla f$ on $\mathcal{T}(\tau_{\mathcal{M}})$ for some manifold $\mathcal{M}$. We will often assume the following approximation condition on a function $s\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$:
$\lVert s(x)-\pi(x)\rVert<\epsilon\,,\quad\lVert s^{\prime}(x)-P_{0}(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (43)
@@ Appendix E Approximate Riemannian gradient flow with landing
$G_{s}^{\eta}(x)=s^{\prime}(x)\nabla f(x)+\eta(x-s(x))$ (44)
E.1 Stationary points of (8) and optimality criteria
@@
$\lVert G_{s}^{\eta}(x_{*})\rVert\leq\delta\,.$ (45)
@@
$\lVert\eta(s(x_{*})-x_{*})-\eta(\pi(x_{*})-x_{*})\rVert\leq\eta\epsilon\,$,
@@
$\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-s^{\prime}(x_{*})\nabla f(s^{\prime}(x_{*}))\rVert$
$\hskip 56.9055pt\quad\leq\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-P_{0}(\pi(x_{*}))\nabla f(s(x_{*}))\rVert$
$\hskip 56.9055pt\qquad+\lVert P_{0}(\pi(x_{*}))\nabla f(s(x_{*}))-P_{0}(x_{*})\nabla f(s(x_{*}))\rVert$
$\hskip 56.9055pt\qquad+\lVert P_{0}(x_{*})\nabla f(s(x_{*}))-s^{\prime}(x_{*})\nabla f(s(x_{*}))\rVert$
@@
where we have used that $P_{0}(\pi(x_{*}))$ is an orthogonal projection and hence has unit norm. By Lemma 23 applied to $z_{1}=s^{\prime}(x_{*})\nabla f(s(x_{*}))$ and $z_{2}=\eta(s(x_{*})-x_{*})$, Lemma 9 and the fact that $P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))=\operatorname{grad}_{\mathcal{M}}f(\pi(x_{*}))$, the inequality (46) follows. ∎
E.2 Lie derivative of manifold distance
@@
$-\left\langle\nabla\operatorname{d}(x),G_{s}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$
@@
$\left\langle x-\pi(x),-G_{s}^{\eta}(x)\right\rangle$
$\hskip 56.9055pt=-2\eta\operatorname{d}(x)+\left\langle\pi(x)-x,(s^{\prime}(x)-P_{0}(x))\nabla f(s(x))\right\rangle$
$\hskip 56.9055pt\qquad+\eta\left\langle x-\pi(x),s(x)-\pi(x)\right\rangle$
$\hskip 56.9055pt\leq-2\eta\operatorname{d}(x)+\lVert\pi(x)-x\rVert\lVert s^{\prime}(x)-P_{0}(x)\rVert\lVert\nabla f(s(x))\rVert$
$\hskip 56.9055pt\qquad+\eta\lVert\pi(x)-x\rVert\lVert s(x)-\pi(x)\rVert$
@@
E.3 Exact landing with $\sigma=0$
@@
Notice that $\{x(t)\mid t\geq 0\}\subseteq\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$ being compact. Since $\mathcal{T}\to\mathbb{R}:z\mapsto\lVert\operatorname{grad}f(z)\rVert_{H_{z}}^{2}$ is continuous, it is uniformly continuous on $\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$. This, together with (47) implies, by Barbalat’s lemma Farkas & Wegner (2016), that $\lim_{t\to\infty}\lVert\operatorname{grad}f(x(t))\rVert_{H_{x(t)}}^{2}=0$. Clearly any accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $H_{x_{*}}^{-1}\operatorname{grad}f(x_{*})=0$, i.e. $\operatorname{grad}f(x_{*})=0$. Then $\pi(x_{*})\in\mathcal{M}$ is an accumulation point of $\{\pi(x(t))\mid t\geq 0\}\subseteq\mathcal{M}$ and
@@
E.4 Proof of Theorem 3
Note first that, by Theorem 1, if (7) hold with $\epsilon\to\epsilon^{\prime}$, then 43 are satisfied with $\epsilon=\epsilon^{\prime}+K\sigma\lvert\log(\sigma)\rvert^{3}$ for $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$. In the following we will write $\epsilon>0$ for the latter quantity. First let us analyze the manifold distance of (8). We have by Lemma 19 whenever $x(t)\in\mathcal{T}(\tau)$ that
$\frac{\operatorname{d}}{\operatorname{d}t}\operatorname{d}(x)=-\left\langle\nabla\operatorname{d}(x),G_{s}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\,$,
@@ E.4 Proof of Theorem 3
Note that $\tau_{0}<\tau$ if $\epsilon<2\tau/(1+C/\eta)$. In particular if $\operatorname{dist}_{\mathcal{M}}(x(0))\in[\tau_{0},\tau)$, then $\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$ is invariant w.r.t. the flow (8) and $x(t)$ exists for all $t\geq 0$. Moreover, by a similar argument as in the standard proof of Lyapunov’s direct method Khalil & Grizzle (2002), for each $\delta>0$ there exists some $T>0$ such that for all $t\geq T$ it holds that $x(t)\in\mathcal{T}(\tau_{0}+\delta)$. If $s$ is a gradient field, i.e. $s=\nabla g$ for some function $g\in C^{1}(\mathbb{R}^{d})$, then $V(x):=f(s(x))+\eta(\lVert x\rVert^{2}/2-g(x))$ satisfies $\nabla V(x)=G_{s}^{\eta}(x)$. Similar as in the proof of Theorem 20 we can take $G_{s}$ as a Lyapunov function to obtain
$\frac{\operatorname{d}}{\operatorname{d}t}V(x)=-\lVert G_{s}^{\eta}(x)\rVert^{2}\,$
along the dynamics (8). By a similar argument $V(x)$ is non-increasing and every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $G_{s}^{\eta}(x)(x_{*})=0$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$. Note that the condition $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\tau)$ in Lemma 18 is satisfied when $\epsilon\leq\tau$. Applying Lemma 18 with $\delta=0$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and proves Theorem 3 for the case when $s$ is a gradient field. Now consider the more general case when $s$ is not necessarily a gradient field. In this case consider $V(x)=f(s(x))+\eta\operatorname{d}(x)$ to obtain
$\frac{\operatorname{d}}{\operatorname{d}t}V(x)$ $=-\left\langle s^{\prime}(x)\nabla f(s(x))+\eta(x-\pi(x)),G_{s}^{\eta}(x)\right\rangle$
$\leq-\lVert G_{s}^{\eta}(x)\rVert^{2}+\eta\lVert s(x)-\pi(x)\rVert\lVert G_{s}^{\eta}(x)\rVert$
$\leq-\lVert G_{s}^{\eta}(x)\rVert^{2}+\eta\epsilon\lVert G_{s}^{\eta}(x)\rVert\,.$
Thus, if $\lVert G_{s}^{\eta}(x)\rVert\geq\eta\epsilon$ we have $\frac{\operatorname{d}}{\operatorname{d}t}V(x)\leq 0$. By a barrier function argument Khalil & Grizzle (2002) this implies that every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $\lVert G_{s}^{\eta}(x_{*})\rVert\leq\eta\epsilon$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$. Again applying Lemma 18 with $\delta=\eta\epsilon$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and finishes the proof.
Appendix F Discretized Riemannian gradient flow and descent

### v2

@@ Appendix D Variance-exploding diffusion models
Appendix E Proof of Theorem 1
E.1 Nonasymptotic Laplace method
@@
for some functions $f,h:\mathcal{V}\to\mathbb{R}$ from an open set $\mathcal{V}\subseteq\mathbb{R}^{k}$, where $h$ is non-negative and attains a unique minimum in, say, $z=0\in\mathcal{V}$. While there have been results providing such an error estimation to the first order expansion Inglot and Majerski (2014); Majerski (2015); Lapinski (2019), we need an error estimate up to the second order expansion, as provided in Theorem 16. For the sake of completeness we also include the statement and proof for the first order expansion in Theorem 15. Before we state the results, we need to introduce some notation and conventions. For any $f\in C^{0}(\mathcal{V})$ and $h\in C^{2}(\mathcal{V})$, respectively we write
@@
E.2 Local version of Theorem 1
@@
E.3 Bounds in terms of distribution and manifold parameters
@@ E.3.1 Conditions (27) and (29) for $h$
E.3.2 Lower bound on $\lambda_{\min}(\Sigma)$ and upper bound on $\sqrt{\det\Sigma}$
@@ E.3.3 Condition (27) for $\overline{f}_{2}^{jl}$
E.3.4 Upper bound on $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}$, $\lVert f_{1}^{j}\rVert_{L^{1}(\mathcal{V})}$ and $\lVert\overline{f}_{2}^{jl}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}$ and $\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert$
@@
E.4 Proof of Theorem 1
We apply Theorem 17 to every point $x\in\mathcal{T}(\tau)$ and $p=\pi(x)$ with the graph chart $\psi=\psi_{p}$. It remains to provide universal constants for the bounds in Theorem 17 (i) and (ii). First we bound $E_{1}(\sigma;f_{0},h,r,\theta)$ with $E_{1}$ given in (34). By Section E.3 we can take $r_{\tau_{\mathcal{M}},M}=\eta=\min\{\tau_{\mathcal{M}},M\}/4$, $c_{\tau_{\mathcal{M}},M}=\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}$ and $C_{\tau_{\mathcal{M}},M,\mu}=\max\{C_{\tau_{\mathcal{M}},M}^{(1)},C_{\tau_{\mathcal{M}},M,\mu}^{(2)}\}$ and hence, due to $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}\leq 1$,
@@ E.4 Proof of Theorem 1
E.5 A useful corollary to Theorem 1
@@
Suppose that for some $\texttt{v}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\tau\in(0,\tau_{\mathcal{M}})$ and $\epsilon\leq\tau$ we have
$\lVert\texttt{v}(x)-\pi(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (45)
Then $\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$. Moreover, for $x\in\mathcal{T}(\epsilon)$ we have $\lVert\texttt{v}(x)-x\rVert\leq 2\epsilon$ and for $x\in\mathcal{T}(\tau)$ we have $\lVert\texttt{v}(x)-x\rVert\leq 2\tau$.
@@
From (45) and the fact that $\pi(x)\in\mathcal{M}$ we clearly have $\texttt{v}(x)\in\mathcal{T}(\epsilon)$ for any $x\in\mathcal{T}(\tau)$, i.e. $\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$. The last claims follows then from the triangle inequality $\lVert\texttt{v}(x)-x\rVert\leq\lVert\texttt{v}(x)-\pi(x)\rVert+\lVert\pi(x)-x\rVert$. ∎
Appendix F Approximate Riemannian gradient flow with landing
In this section $C=\lVert\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})})$ denote the supremum of $\nabla f$ and the Lipschitz constant of $\nabla f$ on $\mathcal{T}(\tau_{\mathcal{M}})$ for some manifold $\mathcal{M}$. We will often assume the following approximation condition on a function $\texttt{v}\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$:
$\lVert\texttt{v}(x)-\pi(x)\rVert<\epsilon\,,\quad\lVert\texttt{v}^{\prime}(x)-P_{0}(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (46)
@@ Appendix F Approximate Riemannian gradient flow with landing
$G_{\texttt{v}}^{\eta}(x)=\texttt{v}^{\prime}(x)\nabla f(x)+\eta(x-\texttt{v}(x))$ (47)
F.1 Stationary points of (8) and optimality criteria
@@
$\lVert G_{\texttt{v}}^{\eta}(x_{*})\rVert\leq\delta\,.$ (48)
@@
$\lVert\eta(\texttt{v}(x_{*})-x_{*})-\eta(\pi(x_{*})-x_{*})\rVert\leq\eta\epsilon\,$,
@@
$\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-\texttt{v}^{\prime}(x_{*})\nabla f(\texttt{v}(x_{*}))\rVert$
$\hskip 56.9055pt\quad\leq\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-P_{0}(\pi(x_{*}))\nabla f(\texttt{v}(x_{*}))\rVert$
$\hskip 56.9055pt\qquad+\lVert P_{0}(\pi(x_{*}))\nabla f(\texttt{v}(x_{*}))-P_{0}(x_{*})\nabla f(\texttt{v}(x_{*}))\rVert$
$\hskip 56.9055pt\qquad+\lVert P_{0}(x_{*})\nabla f(\texttt{v}(x_{*}))-\texttt{v}^{\prime}(x_{*})\nabla f(\texttt{v}(x_{*}))\rVert$
@@
where we have used that $P_{0}(\pi(x_{*}))$ is an orthogonal projection and hence has unit norm. By Lemma 25 applied to $z_{1}=\texttt{v}^{\prime}(x_{*})\nabla f(\texttt{v}(x_{*}))$ and $z_{2}=\eta(\texttt{v}(x_{*})-x_{*})$, Lemma 11 and the fact that $P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))=\operatorname{grad}_{\mathcal{M}}f(\pi(x_{*}))$, the inequality (49) follows. ∎
F.2 Lie derivative of manifold distance
@@
$-\left\langle\nabla\operatorname{d}(x),G_{\texttt{v}}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$
@@
$\left\langle x-\pi(x),-G_{\texttt{v}}^{\eta}(x)\right\rangle$
$\hskip 56.9055pt=-2\eta\operatorname{d}(x)+\left\langle\pi(x)-x,(\texttt{v}^{\prime}(x)-P_{0}(x))\nabla f(\texttt{v}(x))\right\rangle$
$\hskip 56.9055pt\qquad+\eta\left\langle x-\pi(x),\texttt{v}(x)-\pi(x)\right\rangle$
$\hskip 56.9055pt\leq-2\eta\operatorname{d}(x)+\lVert\pi(x)-x\rVert\lVert\texttt{v}^{\prime}(x)-P_{0}(x)\rVert\lVert\nabla f(\texttt{v}(x))\rVert$
$\hskip 56.9055pt\qquad+\eta\lVert\pi(x)-x\rVert\lVert\texttt{v}(x)-\pi(x)\rVert$
@@
F.3 Exact landing with $\sigma=0$
@@
Notice that $\{x(t)\mid t\geq 0\}\subseteq\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$ being compact. Since $\mathcal{T}\to\mathbb{R}:z\mapsto\lVert\operatorname{grad}f(z)\rVert_{H_{z}}^{2}$ is continuous, it is uniformly continuous on $\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$. This, together with (50) implies, by Barbalat’s lemma Farkas and Wegner (2016), that $\lim_{t\to\infty}\lVert\operatorname{grad}f(x(t))\rVert_{H_{x(t)}}^{2}=0$. Clearly any accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $H_{x_{*}}^{-1}\operatorname{grad}f(x_{*})=0$, i.e. $\operatorname{grad}f(x_{*})=0$. Then $\pi(x_{*})\in\mathcal{M}$ is an accumulation point of $\{\pi(x(t))\mid t\geq 0\}\subseteq\mathcal{M}$ and
@@
F.4 Proof of Theorem 3
Note first that, by Theorem 1, if (7) hold with $\epsilon\to\epsilon^{\prime}$, then 46 are satisfied with $\epsilon=\epsilon^{\prime}+K\sigma\lvert\log(\sigma)\rvert^{3}$ for $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$. In the following we will write $\epsilon>0$ for the latter quantity. First let us analyze the manifold distance of (8). We have by Lemma 21 whenever $x(t)\in\mathcal{T}(\tau)$ that
$\frac{\operatorname{d}}{\operatorname{d}t}\operatorname{d}(x)=-\left\langle\nabla\operatorname{d}(x),G_{\texttt{v}}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\,$,
@@ F.4 Proof of Theorem 3
Note that $\tau_{0}<\tau$ if $\epsilon<2\tau/(1+C/\eta)$. In particular if $\operatorname{dist}_{\mathcal{M}}(x(0))\in[\tau_{0},\tau)$, then $\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$ is invariant w.r.t. the flow (8) and $x(t)$ exists for all $t\geq 0$. Moreover, by a similar argument as in the standard proof of Lyapunov’s direct method Khalil and Grizzle (2002), for each $\delta>0$ there exists some $T>0$ such that for all $t\geq T$ it holds that $x(t)\in\mathcal{T}(\tau_{0}+\delta)$. If v is a gradient field, i.e. $\texttt{v}=\nabla g$ for some function $g\in C^{1}(\mathbb{R}^{d})$, then $V(x):=f(\texttt{v}(x))+\eta(\lVert x\rVert^{2}/2-g(x))$ satisfies $\nabla V(x)=G_{\texttt{v}}^{\eta}(x)$. Similar as in the proof of Theorem 22 we can take $G_{\texttt{v}}$ as a Lyapunov function to obtain
$\frac{\operatorname{d}}{\operatorname{d}t}V(x)=-\lVert G_{\texttt{v}}^{\eta}(x)\rVert^{2}\,$
along the dynamics (8). By a similar argument $V(x)$ is non-increasing and every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $G_{\texttt{v}}^{\eta}(x)(x_{*})=0$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$. Note that the condition $\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\tau)$ in Lemma 20 is satisfied when $\epsilon\leq\tau$. Applying Lemma 20 with $\delta=0$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and proves Theorem 3 for the case when $s$ is a gradient field. Now consider the more general case when v is not necessarily a gradient field. In this case consider $V(x)=f(\texttt{v}(x))+\eta\operatorname{d}(x)$ to obtain
$\frac{\operatorname{d}}{\operatorname{d}t}V(x)$ $=-\left\langle\texttt{v}^{\prime}(x)\nabla f(\texttt{v}(x))+\eta(x-\pi(x)),G_{\texttt{v}}^{\eta}(x)\right\rangle$
$\leq-\lVert G_{\texttt{v}}^{\eta}(x)\rVert^{2}+\eta\lVert\texttt{v}(x)-\pi(x)\rVert\lVert G_{\texttt{v}}^{\eta}(x)\rVert$
$\leq-\lVert G_{\texttt{v}}^{\eta}(x)\rVert^{2}+\eta\epsilon\lVert G_{\texttt{v}}^{\eta}(x)\rVert\,.$
Thus, if $\lVert G_{\texttt{v}}^{\eta}(x)\rVert\geq\eta\epsilon$ we have $\frac{\operatorname{d}}{\operatorname{d}t}V(x)\leq 0$. By a barrier function argument Khalil and Grizzle (2002) this implies that every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $\lVert G_{\texttt{v}}^{\eta}(x_{*})\rVert\leq\eta\epsilon$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$. Again applying Lemma 20 with $\delta=\eta\epsilon$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and finishes the proof.
Appendix G Discretized Riemannian gradient flow and descent